OpenStudy (anonymous):

write an equation of the line that passes through points (3,7) and (5,11)

OpenStudy (anonymous):

A. y-7=5(x-3) B. y-7=2(x-3) C. y+7=5(x-3) D. y+7=2(x+3)

OpenStudy (uri):

First find slope. \[m= \frac{ (y2-y1)}{ (x2-x1) }\]

OpenStudy (uri):